Circular No. 8381

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2004dn IN UGC 2069 Further to IAUC 8377, J. Graham and W. Li report the LOSS discovery

of an apparent supernova on an unfiltered KAIT image taken on Aug. 2.48 UT (mag ∼ 17.5). The new object was confirmed on an earlier image taken on July 29.42 (mag ∼ 18.5). SN 2004dn is located at α = 2^{h}35^{m}37^{s}.30, δ = +37^{o}37′54′′.2 (equinox 2000.0), which is 0′′.5 west and 26′′.1 south of the nucleus of UGC 2069. A KAIT image taken on Jan. 23.21 showed nothing at this position (limiting mag ∼ 19.5). SN 1961P was located 36′′ south of the galaxy’s nucleus.

F. Patat and G. Pignata, European Southern Observatory; and S. Benetti, Istituto Nazionale di Astrofisica, on behalf of the ERTN (see IAUC 7954 and 8344 — not IAUC 8341, as cited incorrectly on IAUC 8378, 8379), report that a fully reduced spectrum (range 320–900 nm; resolution 1.6 nm) of SN 2004dn, obtained on Aug. 5.15 UT by J. Aceituno at the Calar Alto Observatory 2.2-m telescope (+ CAFOS), shows it to be a type-Ic supernova. The spectrum is very similar to those of the type-Ic SNe 1991A and 1994ai (Asiago-ESO supernova archive, unpublished; IAUC 5155, 6120), a few days before maximum light. Photospheric expansion velocities, deduced from the absorption minima of Si II 635.5-nm and O I 777.4-nm, are ∼ 9600 km/s and 10700 km/s, respectively. The recession velocity of UGC 2069 at the supernova position, measured from the unresolved Hα and Hβ arising in the underlying H II region, is ∼ 3950 km/s.

NOVA SCORPII 2004 No. 2 H. Yamaoka, Kyushu University, forwards the following position for

this object (cf. IAUC 8380) measured by K. Itagaki from an unfiltered CCD image taken on Aug. 5.475 UT (with poor sky conditions; the new object was at mag 8.0): α = 17^{h}29^{m}18^{s}.81, δ = −31^{o}46′01′′.5 (equinox 2000.0). This appears nearly identical with the location of a very red catalogued star whose USNO-A2.0 position end figures are 18^{s}.818, 01′′.68 (blue mag 19.7, red mag 17.4), and whose 2MASS position end figures are 18^{s}.82, 01′′.5 (J = 14.4). Yamaoka adds that a low-resolution spectrum of the new object, taken by M. Fujii, Bisei, Okayama, Japan, on Aug. 5.49, shows broad Hα and Hβ emissions with P-Cyg profiles (Hα FWHM = 3000 km/s), along with several Fe II and O I emission lines — suggesting that the object is a classical nova. W. Liller, Viña del Mar, Chile, writes that the nova was fainter than mag 11.0 on a Technical Pan film exposed (with an orange filter) on July 19.40.

2004 August 5  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2004 CBAT  Daniel W. E. Green